sealake101:

PLZ help me... The tuition costs, C, for a local community college are modeled by C(h) = 250 + 200h, where h represents the number of credit hours taken. The local state university has tuition costs, S, modeled by the function S(h) = 300 + 180h. How many credit hours will a student have to take for the two tuition costs to be equal? Round the answer to the nearest tenth of an hour. 1. 250 + 200h = 300 + 180h 2. 250 + 200h = 300 + 180h over − 180h − 180h 3. 250 + 20h = 300 h=____________ credit hours

supie:

\[ C (h) = S (h)\]\[250 + 200h = 300 + 180h\]\[200h – 180h = 300 – 250\]\[20h = 50\] Does that help?

sealake101:

yeah, thx

supie:

Ok Np.

Adverturetime57:

1. 250 + 200h = 300 + 180h 2. 250 + 200h = 300 + 180h over − 180h − 180h 3. 250 + 20h = 300 so 250+20 =70 so i belive you round 70 to the nearest 10 hope this helps

sealake101:

ok thank u!